OpenStudy (anonymous):

What is 14x+3=19x+23

pooja195 (pooja195):

Subtract 19x from both sides. 14x+3−19x=19x+23−19x −5x+3=23

pooja195 (pooja195):

Subtract 3 from both sides. −5x+3−3=23−3 −5x=20 Divide both sides my -5 x=20/-5

OpenStudy (anonymous):

Divide both sides by -5*

OpenStudy (anonymous):

I ÷20 and -5